OpenStudy (anonymous):

find two integers whose sum is 96 and whose product is 1728

OpenStudy (anonymous):

a + b = 96 ab = 1728 Note (x-a)(x-b) = x^2 - (a+b)x + ab So we can say the solutions are the solutions to the quadratic: x^2 - 96x+1728 = 0

OpenStudy (anonymous):

y do we take the formula (x-a)(x-b)

OpenStudy (anonymous):

Because I'm a badass. It's just an (in my opinion) slightly slicker operation. You can go a = 96 - b => ab = (96-b)b = 1728 etc, and it leads to the same thing.

OpenStudy (anonymous):

Then plug that back in for a.

OpenStudy (anonymous):

okay thanx